# PINSAT: Parallelized Interleaving of Graph Search and Trajectory Optimization for Kinodynamic Motion Planning

### 17 Jan 2024 v1

@@ 5.1 B-Spline Optimization
Transcription for Optimizing B-splines
The derivative of the B-spline curve $\boldsymbol{\phi}^{(j)}(t)$ in terms of the derivative of the B-spline basis $N_{i,k}^{(j)}(t)$ is given in Eq. 8. Using this relation the derivative of the B-spline curve with the trajectory coordinate $\boldsymbol{\psi}^{(j)}(u)$ in terms of the derivative’s control points $\textbf{p}_{i}^{(j)}$ can be derived as (Piegl and Tiller 1996)
where $\textbf{p}_{i}^{(j)}$ is given as
Now we can apply the decoupling introduced above to the optimization in Eq. 3 with the new trajectory coordinate $u$ and the map $\boldsymbol{\psi}$ introduced above. Given two boundary states $\mathbf{x}^{\prime},\mathbf{x}^{\prime\prime}$, the optimization in Eq. 3 can be transcribed into a parameter optimization problem over the control points of the B-spline trajectory $\textbf{P}=\{\textbf{p}_{0},\ldots,\textbf{p}_{m}\}$ and the duration of the trajectory $T$ using Eq. 14 as
s.t. $0<t_{\text{min}}\leq T\leq t_{\text{max}}$ (15b)
$\textbf{p}_{0}=\mathbf{x}^{S},\textbf{p}_{m}=\mathbf{x}^{G}\in\mathcal{G}$ (15c)
where the constraint Eq. 15d comes from Eq. 14 and $\boldsymbol{\psi}_{\textbf{P}},T$ is the trajectory reconstructed during every iteration using P and $T$ as explained below. From the properties of B-splines (De Boor 1978), we know that the B-spline curve is entirely contained inside the convex hull of its control points. We leverage this property to formulate Eq. 15a and Eq. 15d. So it is enough to just enforce the trajectory derivative constraints at the control points (Eq. 15d) instead of checking if the curve satisfies them for all $u$ (and therefore for all $t$).
Remark 3.
The constraint on the control points in Eq. 15d is a sufficient but not necessary condition to enforce the limits on derivatives. In other words, it is possible to have control points outside the limits ±$\pm T^{j}\boldsymbol{\phi}^{(j)}_{\text{lim}}$ and still have the curve satisfy these limits.
Note that Eq. 5.1 is a nonlinear program (NLP) for $j>1$ as it has quadratic (or cubic, quintic, etc) constraints in the decision variable $T$ (Eq. 15d). However, the following relaxation obtained by fixing the duration to $t_{\text{max}}$ is convex if $\mathcal{X}^{\text{free}}$ is convex
B-spline Trajectory Reconstruction
Once the control point vector P and the duration of trajectory $T$ are found by solving Eq. 5.1, the trajectory and its derivatives can be computed given the choice of B-spline basis $N_{i,k}(t)$ (Eq. 2) and the knot vector T (Eq. 4) using Eq. 6 and Eq. 12. We also use this reconstruction to validate Eq. 15e within every iteration of the solver routine for obstacle avoidance, early termination and constraint satisfaction.
@@
15:        if such an edge does not exist then
16:           UNLOCK
17:           wait until OPEN or BE change
18:           LOCK
19:           continue
20:        end if
21:        if $\mathbf{x}\in\mathcal{G}$ then
22:           $terminate=\text{True}$
23:           UNLOCK
24:           return  $\mathbf{x}.traj$
25:        else
26:           UNLOCK
27:           while $(\mathbf{x},\mathbf{a})$ has not been assigned a thread do
28:              for $i=1:N_{t}$ do
29:                 if thread $i$ is available then
30:                    if thread $i$ has not been spawned then
@@
32:                    end if
33:                    Assign $(\mathbf{x},\mathbf{a})$ to thread $i$
34:                 end if
35:              end for
36:           end while
37:           LOCK
38:        end if
39:     end while
40:     $terminate=\text{True}$
41:     UNLOCK
42:  end procedure
5.2 Low Dimensional Graph Search
The low-D search runs w-eA* (Mukherjee, Aine, and Likhachev 2022a) with variations that enable its interleaving with the full-D trajectory optimization. Using w-eA* instead of wA* for the low-D search provides a systematic framework for parallelizing the expensive trajectory optimization via its parallelized variant w-ePA*SE. This is because, unlike in wA*, where the basic operation of the search loop is the expansion of a state, in w-eA*, the basic operation is the expansion of an edge.
@@ 5.2 Low Dimensional Graph Search
5:        end if
6:     end while
7:  end procedure
8:  procedure Expand($(\mathbf{x},\mathbf{a})$)
9:     LOCK
10:     if $\mathbf{a}=\mathbf{a^{d}}$ then
11:        insert $\mathbf{x}$ in BE
12:        for $\mathbf{a}\in\mathcal{A}$ do
13:           $f\left((\mathbf{x},\mathbf{a})\right)=g(\mathbf{x})+h(\mathbf{x})$
14:           insert $(\mathbf{x},\mathbf{a})$ in OPEN with $f\left((\mathbf{x},\mathbf{a})\right)$
15:        end for
16:     else
17:        UNLOCK
18:        $\mathbf{x}^{\prime},c\left((\mathbf{x},\mathbf{a})\right)\leftarrow\textsc{GenerateSuccessor}\left((\mathbf{x},\mathbf{a})\right)$
19:        LOCK
@@
21:           $\boldsymbol{\phi}(\mathbf{x}^{S}\mathbf{x}^{\prime})=\textsc{GenerateTrajectory}(\mathbf{x},\mathbf{x}^{\prime})$
22:           if $g(\mathbf{x}^{\prime})>c(\boldsymbol{\phi}(\mathbf{x}^{S}\mathbf{x}^{\prime}))$ then
23:              $g(\mathbf{x}^{\prime})=c(\boldsymbol{\phi}(\mathbf{x}^{S}\mathbf{x}^{\prime}))$
24:              $\mathbf{x}^{\prime}.parent=\mathbf{x}$
25:              $\mathbf{x}^{\prime}.traj=\boldsymbol{\phi}_{\mathbf{x}^{S}\mathbf{x}^{\prime}}$
26:              $f\left((\mathbf{x}^{\prime},\mathbf{a^{d}})\right)=g(\mathbf{x}^{\prime})+h(\mathbf{x}^{\prime})$
@@
27:              insert/update $(\mathbf{x}^{\prime},\mathbf{a^{d}})$ in OPEN with $f\left((\mathbf{x}^{\prime},\mathbf{a^{d}})\right)$
28:           end if
29:        end if
30:        $n\_successors\_generated(\mathbf{x})+=1$
31:        if $n\_successors\_generated(\mathbf{x})=|\mathcal{A}|$ then
32:           remove $\mathbf{x}$ from BE
33:           insert $\mathbf{x}$ in CLOSED
34:        end if
35:     end if
36:     UNLOCK
37:  end procedure
w-eA*
5.3 PINSAT: Parallelized Interleaving of Search And Trajectory Optimization
The pseudocode of PINSAT is given in Alg. 1 along with a graphical illustration in Fig. 2. In graph search terminology, edge evaluation refers to computing the cost and representation of an edge that is part of the underlying graph. The optimizing step of rewiring to a better parent node, common in many search algorithms, is restricted to the edges within the graph. In contrast, a key distinction in PINSAT is that it dynamically generates edges between nodes that are not part of the low-D graph when searching for a full-D trajectory to a node. INSAT searches for this dynamic edge over the set of ancestor nodes leading to the node picked for expansion. Consequently, to make w-eA* and ePA*SE compatible with PINSAT, we search over the set of ancestor edges leading up to the edge picked for expansion (Alg. 3, line 2).
Each time the search expands a state, it removes that state from OPEN and generates the successors as per the discretization. For every low-dimensional successor, we solve a trajectory optimization problem described in the previous sections, to find a corresponding full-dimensional trajectory from the start to the successor. The trajectory optimization has access to the collision checker (implemented through callback mechanisms) and runs until convergence or collision occurrence, whichever comes first. During every iteration of the optimization, the iterates are reconstructed and checked for collision. In the case of collision, if the most recent iterate satisfies all the constraints then it is returned as a solution. If the optimized trajectory is in collision or infeasible, the algorithm enters the repair phase.
Parallelization
@@ 5.3 PINSAT: Parallelized Interleaving of Search And Trajectory Optimization
In w-ePA*SE, to maintain bounded sub-optimality, an edge can only be expanded if it is independent of all edges ahead of it in OPEN and the edges currently being expanded i.e. in set BE (Mukherjee, Aine, and Likhachev 2022a). However, since INSAT and therefore PINSAT are not bounded sub-optimal algorithms, the expensive independence check can be eliminated. Additionally, this increases the amount of parallelization that can be achieved. Once an edge is popped from OPEN 14, it is delegated for expansion to a spawned edge expansion thread. If all of the spawned edges are busy, a new thread is spawned as long as the total number of threads does not exceed the thread budget $N_{t}$ 31. When $N_{t}$ is higher than the number of edges available for expansion at any point in time, only a subset of available threads get spawned. This prevents performance degradation due to the operating system overhead of managing unused threads.
@@
3:        $\boldsymbol{\phi}_{\mathbf{x}^{\prime\prime}\mathbf{x}^{\prime}}=\textsc{Optimize}(\mathbf{x}^{\prime\prime},\mathbf{x}^{\prime})$ Eq. 5.1
@@
5:           $\boldsymbol{\phi}_{\mathbf{x}^{S}\mathbf{x}^{\prime}}=\textsc{WarmOptimize}(\boldsymbol{\phi}_{\mathbf{x}^{S}\mathbf{x}^{\prime\prime}},\boldsymbol{\phi}_{\mathbf{x}^{\prime\prime}\mathbf{x}^{\prime}})$ Eq. 5.1
@@
8:           if $\boldsymbol{\phi}_{\mathbf{x}^{\prime\prime}\mathbf{x}^{\prime}}^{t_{\text{max}}}$ exists then # Solve Eq. 5.1
9:              $\boldsymbol{\phi}_{\mathbf{x}^{\prime\prime}\mathbf{x}^{\prime}}=\textsc{Optimize}(\mathbf{x}^{\prime\prime},\mathbf{x}^{\prime})$ Eq. 5.1 with $k_{\text{min}}$
10:              $\boldsymbol{\phi}_{\mathbf{x}^{S}\mathbf{x}^{\prime}}=\textsc{WarmOptimize}(\boldsymbol{\phi}_{\mathbf{x}^{S}\mathbf{x}^{\prime\prime}},\boldsymbol{\phi}_{\mathbf{x}^{\prime\prime}\mathbf{x}^{\prime}})$ 5.1
@@
12:           end if
13:        end if
14:        return  NULL
15:     end for
16:  end procedure
5.4 Theoretical Analysis
In this section, we will show that PINSAT is a provably complete algorithm, i.e. it finds a solution if one exists. We will begin with a few preliminaries required to prove completeness.
@@ 5.4 Theoretical Analysis
We assume that there exists at least a path on the low-D graph $G$ form $\mathbf{x}^{S}$ to $\mathbf{x}^{G}$ such that its full-D tunnel contains $\boldsymbol{\phi}_{\textbf{x}^{S}\textbf{x}^{G}}$, i.e. $\exists\boldsymbol{\theta}_{\textbf{x}^{S}\textbf{x}^{G}}\in G\mid\exists\boldsymbol{\phi}_{\textbf{x}^{S}\textbf{x}^{G}}\in\tau(\boldsymbol{\theta}_{\textbf{x}^{S}\textbf{x}^{G}})$.
@@
Minimum Order of B-spline:
@@
The tunnel constraints in Eq 17c are transcribed into constraints on the control points similar to how corridor constraints are represented in (Rousseau et al. 2019). In this work, the outer maximization is carried over actions set $\mathcal{A}$, which is typically a significantly smaller set than $\mathcal{E}$. This can be done if the edges in the graph can be grouped as a finite number of action primitives sharing the same tunnel representation. Note that the inner optimization in the above formulation is a linear program and is guaranteed to have at least one solution as long as the feasible region of the decision variable is nonempty and bounded.
@@
The feasible region depends on the limits of the system $\boldsymbol{\phi}^{(j)}_{\text{lim}}$ and the choice of the tunnel $\tau$ which is domain-specific. However, in most cases, a straightforward design of the tunnel (Rousseau et al. 2019) can ensure a nonempty feasible region and a reasonably low $k_{\text{min}}$.
@@
If $\tau(e)$ is convex, then a $\boldsymbol{\phi}_{\mathbf{x}^{\prime}\mathbf{x}^{\prime\prime}}^{\sim}(t)\in\tau(e),\forall t\in[0,t_{\text{max}}]$ that solves Eq. 5.1 to global optimality can be found.
@@
Setting $\mathcal{X}^{\text{free}}=\tau(e)$ make Eq. 16d a convex domain constraint on the decision variables and Eq. 5.1 a convex program. Assuming $t_{\text{max}}$ is large enough, we know that a convex optimization can be solved to global optimality. ∎
@@
If $\boldsymbol{\hat{\phi}}_{\mathbf{x}^{\prime}\mathbf{x}^{\prime\prime}}\in\tau_{L}(e)$, then there exists a $k_{\text{min}}$-th order $\boldsymbol{\phi}_{\mathbf{x}^{\prime}\mathbf{x}^{\prime\prime}}^{*}\in\tau(e)$ that satisfies Eq. 5.1.
@@
It directly follows from the constraint satisfaction for finding $k_{\text{min}}$ in Eq. 5.4. This $k_{\text{min}}$ can be used to choose the B-spline basis for optimizing Eq. 5.1 simultaneously over path length and time to find $\boldsymbol{\phi}_{\mathbf{x}^{\prime}\mathbf{x}^{\prime\prime}}^{*}\in\tau(e)$. ∎
@@
Completeness of PINSAT: If Assumption 1 holds, then PINSAT with $k_{\text{min}}$-th order B-spline optimization is guaranteed to find a $\boldsymbol{\phi}_{\mathbf{x}^{S}\mathbf{x}^{G}}$ that satisfies Eq. 3.
6 Evaluation
pBiRRT +TrajOpt w-ePA*SE +TrajOpt INSAT PINSAT
Threads 5 10 50 120 5 10 50 120 1 5 10 50 120
[…]
Table 1: Mean and standard deviation of planning time and cost for PINSAT and the baselines for different thread budgets.
Figure 3: PINSAT used accelerate preprocessing of a motion library for rapid lookup. Here an ABB arm with a shield attached to its end-effector executes a motion generated by PINSAT to block a ball thrown at it.

### 16 Mar 2024 v2

@@ V-B1 Decoupling Optimization over Trajectory and its Duration
14:        if such an edge does not exist then
15:           UNLOCK
16:           wait until OPEN or BE change
17:           LOCK
18:           continue
19:        if $\mathbf{x}\in\mathcal{G}$ then
20:           $terminate=\text{True}$
21:           UNLOCK
22:           return  $\mathbf{x}.traj$
23:        else
24:           UNLOCK
25:           while $(\mathbf{x},\mathbf{a})$ has not been assigned a thread do
26:              for $i=1:N_{t}$ do
27:                 if thread $i$ is available then
28:                    if thread $i$ has not been spawned then
@@
30:                    Assign $(\mathbf{x},\mathbf{a})$ to thread $i$
31:           LOCK
32:     $terminate=\text{True}$
33:     UNLOCK
V-B2 Transcription for Optimizing B-splines
The derivative of the B-spline curve $\boldsymbol{\phi}^{(j)}(t)$ in terms of the derivative of the B-spline basis $N_{i,k}^{(j)}(t)$ is given in Eq. 6. Using this relation the derivative of the B-spline curve with the trajectory coordinate $\boldsymbol{\psi}^{(j)}(u)$ in terms of the derivative’s control points $\textbf{p}_{i}^{(j)}$ can be derived as [43]
where $\textbf{p}_{i}^{(j)}$ is given as
@@ V-B2 Transcription for Optimizing B-splines
5:  procedure Expand($(\mathbf{x},\mathbf{a})$)
6:     LOCK
7:     if $\mathbf{a}=\mathbf{a^{d}}$ then
8:        insert $\mathbf{x}$ in BE
9:        for $\mathbf{a}\in\mathcal{A}$ do
10:           $f\left((\mathbf{x},\mathbf{a})\right)=g(\mathbf{x})+h(\mathbf{x})$
11:           insert $(\mathbf{x},\mathbf{a})$ in OPEN with $f\left((\mathbf{x},\mathbf{a})\right)$
12:     else
13:        UNLOCK
14:        $\mathbf{x}^{\prime},c\left((\mathbf{x},\mathbf{a})\right)\leftarrow\textsc{GenerateSuccessor}\left((\mathbf{x},\mathbf{a})\right)$
15:        LOCK
@@
17:           $\boldsymbol{\phi}_{\mathbf{x}^{S}\mathbf{x}^{\prime}}=\textsc{GenerateTrajectory}(\mathbf{x},\mathbf{x}^{\prime})$
18:           if $g(\mathbf{x}^{\prime})>c(\boldsymbol{\phi}(\mathbf{x}^{S}\mathbf{x}^{\prime}))$ then
19:              $g(\mathbf{x}^{\prime})=c(\boldsymbol{\phi}(\mathbf{x}^{S}\mathbf{x}^{\prime}))$
20:              $\mathbf{x}^{\prime}.parent=\mathbf{x}$
21:              $\mathbf{x}^{\prime}.traj=\boldsymbol{\phi}_{\mathbf{x}^{S}\mathbf{x}^{\prime}}$
22:              $f\left((\mathbf{x}^{\prime},\mathbf{a^{d}})\right)=g(\mathbf{x}^{\prime})+h(\mathbf{x}^{\prime})$
@@
23:              insert/update $(\mathbf{x}^{\prime},\mathbf{a^{d}})$ in OPEN with $f\left((\mathbf{x}^{\prime},\mathbf{a^{d}})\right)$
24:        $n\_successors\_generated(\mathbf{x})+=1$
25:        if $n\_successors\_generated(\mathbf{x})=|\mathcal{A}|$ then
26:           remove $\mathbf{x}$ from BE
27:           insert $\mathbf{x}$ in CLOSED
28:     UNLOCK
Now we can apply the above decoupling to the optimization in Eq. III with the new trajectory coordinate $u$ and the map $\boldsymbol{\psi}$. Given two boundary states $\mathbf{x}^{\prime},\mathbf{x}^{\prime\prime}$, Eq. III can be transcribed into a parameter optimization problem over the control points of the B-spline trajectory $\textbf{P}=\{\textbf{p}_{0},\ldots,\textbf{p}_{m}\}$ and the duration of the trajectory $T$ using Eq. 12 as
s.t. $0<t_{\text{min}}\leq T\leq t_{\text{max}}$ (13b)
$\textbf{p}_{0}=\mathbf{x}^{\prime},\textbf{p}_{m}=\mathbf{x}^{\prime\prime}$ (13c)
where the constraint Eq. 13d comes from Eq. 12 and $\boldsymbol{\psi}_{\textbf{P},T}$ is the trajectory reconstructed during every iteration using P and $T$ as explained below. From the properties of B-splines [44], we know that the B-spline curve is entirely contained inside the convex hull of its control points. We leverage this property to formulate Eq. 13a and Eq. 13d. So it is enough to just enforce the trajectory derivative constraints at the control points (Eq. 13d) instead of checking if the curve satisfies them for all $u$ (and therefore for all $t$).
Remark 3.
The constraint on the control points in Eq. 13d is a sufficient but not necessary to enforce the limits on derivatives. In other words, it is possible to have control points outside the limits ±$\pm T^{j}\boldsymbol{\phi}^{(j)}_{\text{lim}}$ and still have the curve satisfy these limits.
Note that Eq. V-B2 is a nonlinear program (NLP) for $j>1$ as it has quadratic (or cubic, quintic, etc) constraints in the decision variable $T$ (Eq. 13d). However, the following relaxation obtained by fixing the duration to $t_{\text{max}}$ is convex if $\mathcal{X}^{\text{free}}$ is convex
V-B3 B-spline Trajectory Reconstruction
Once the control point vector P and the duration of trajectory $T$ are found by solving Eq. V-B2, the trajectory and its derivatives can be computed given the choice of B-spline basis $N_{i,k}(t)$ (Eq. 2) and the knot vector T (Eq. 3a) using Eq. 4 and Eq. 10. We also use this reconstruction to validate Eq. 13e within every iteration of the solver routine for obstacle avoidance, early termination and constraint satisfaction.
V-C PINSAT: Parallelized Interleaving of Search And Trajectory Optimization
@@
The pseudocode of PINSAT is given in Alg. 1 along with a graphical illustration in Fig. 2. The caption in Fig. 2 is intentionally detailed and not repeated in this section. In w-ePA*SE, to maintain bounded suboptimality, an edge can only be expanded if it is independent of all edges ahead of it in OPEN and the edges currently being expanded, i.e. in set BE (Alg. 1, line 9[8]. However, since INSAT and, therefore, PINSAT are not bounded suboptimal algorithms, the expensive independence check can be eliminated. Additionally, this increases the amount of parallelization that can be achieved. Alg. 1 begins by initializing g-values and inserting the start state with the dummy edge ($\mathbf{x}^{S},\mathbf{a^{d}}$) in OPEN (line 4-6). Once an edge is popped from OPEN (Alg. 1, line 13), it is delegated for expansion to a spawned edge expansion thread. If all of the spawned edges are busy, a new thread is spawned as long as the total number of threads does not exceed the thread budget $N_{t}$ (Alg. 1, line  29). When $N_{t}$ is higher than the number of edges available for expansion at any point in time, only a subset of available threads are spawned. This prevents performance degradation due to the operating system overhead of managing unused threads.
The pseudocode for asynchronous edge expansion is given in Alg. 2. If the edge selected for expansion is a dummy edge, then it is first replaced by unevaluated yet real edges in OPEN using the low-D action set $\mathcal{A}$ (line 7-11) and the priority value of its source state. Otherwise, the successor edge is generated and the algorithm enters the trajectory optimization routine (line 12-27) if the edge is neither in CLOSED nor in BE (line 16).
The trajectory optimization routine (provided in Alg. 3) receives the set of ancestors of the edge being expanded (line 2). We solve the optimization problem described in the previous sections (Eq. V-B2) to find a corresponding full-dimensional trajectory from the start to the successor. This optimization is carried out in two steps. Only the trajectory generation between the ancestor node and the successor node, which is typically very short-horizon, must be optimized from scratch (line 3, 9). Finding the entire trajectory to the successor from the start state can be warm-started (line 5, 10) using the incoming trajectory to the ancestor. This is possible as a result of the dynamic programming in the low-D search which guarantees that an edge out of node which does not contain a valid incoming trajectory will never be expanded. In heuristic search, edge evaluation refers to computing the cost and representation of an edge that is part of the underlying graph. The optimization step of rewiring to a better parent node, common in many search algorithms, is restricted to the edges within the graph. In contrast, a key distinction in INSAT and therefore PINSAT is that it optimizes for trajectories between nodes not connected by an edge in the low-D graph when searching for a full-D trajectory to a node. INSAT generates this runtime edge by searching over the set of ancestor nodes that lead to the node picked for expansion. Consequently, to make w-eA* and ePA*SE compatible with PINSAT, we search the set of ancestor edges leading to the edge chosen for expansion. Trajectory optimization has access to the collision checker (implemented through callback mechanisms) and runs until convergence or collision occurrence, whichever comes first. During every iteration of the optimization, the iterates are reconstructed and checked for collision. In the case of collision, if the most recent iterate satisfies all the constraints, then it is returned as a solution. If the optimized trajectory is invalid, the algorithm moves to the next ancestor (line 2).
@@
3:        $\boldsymbol{\phi}_{\mathbf{x}^{\prime\prime}\mathbf{x}^{\prime}}=\textsc{Optimize}(\mathbf{x}^{\prime\prime},\mathbf{x}^{\prime})$ Eq. V-B2
@@
5:           $\boldsymbol{\phi}_{\mathbf{x}^{S}\mathbf{x}^{\prime}}=\textsc{WarmOptimize}(\boldsymbol{\phi}_{\mathbf{x}^{S}\mathbf{x}^{\prime\prime}},\boldsymbol{\phi}_{\mathbf{x}^{\prime\prime}\mathbf{x}^{\prime}})$ Eq. V-B2
@@
8:           if $\boldsymbol{\phi}_{\mathbf{x}^{\prime\prime}\mathbf{x}^{\prime}}^{t_{\text{max}}}$ exists then # Solve Eq. V-B2
9:              $\boldsymbol{\phi}_{\mathbf{x}^{\prime\prime}\mathbf{x}^{\prime}}=\textsc{Optimize}(\mathbf{x}^{\prime\prime},\mathbf{x}^{\prime})$ Eq. V-B2 with $k_{\text{min}}$
10:              $\boldsymbol{\phi}_{\mathbf{x}^{S}\mathbf{x}^{\prime}}=\textsc{WarmOptimize}(\boldsymbol{\phi}_{\mathbf{x}^{S}\mathbf{x}^{\prime\prime}},\boldsymbol{\phi}_{\mathbf{x}^{\prime\prime}\mathbf{x}^{\prime}})$ V-B2
@@
12:        return  NULL
V-D Theoretical Analysis
In this section, we will show that PINSAT is a provably complete algorithm, i.e. it finds a solution if one exists. We will begin with some preliminaries required to prove completeness.
@@ V-D Theoretical Analysis
We assume that there exists at least a path on the low-D graph $G$ from $\mathbf{x}^{S}$ to $\mathbf{x}^{G}$ such that its full-D tunnel contains $\boldsymbol{\phi}_{\textbf{x}^{S}\textbf{x}^{G}}$, i.e. $\exists\boldsymbol{\theta}_{\textbf{x}^{S}\textbf{x}^{G}}\in G\mid\exists\boldsymbol{\phi}_{\textbf{x}^{S}\textbf{x}^{G}}\in\tau(\boldsymbol{\theta}_{\textbf{x}^{S}\textbf{x}^{G}})$.
@@
Figure 3: PINSAT used to accelerate preprocessing of a motion library for rapid lookup. Here, an ABB arm with a shield attached to its end-effector executes a motion generated by PINSAT to block a ball launched toward it.
pBiRRT +TrajOpt w-ePA*SE +TrajOpt INSAT PINSAT
Threads 5 10 50 120 5 10 50 120 1 5 10 50 120
[…]
TABLE I: Mean and standard deviation of planning time and cost for PINSAT and the baselines for different thread budgets.
V-D1 Minimum Order of B-spline
@@
The tunnel constraints in Eq 15c are transcribed into constraints on the control points similar to how corridor constraints are represented in [45]. In this work, the outer maximization is carried over actions set $\mathcal{A}$, which is typically a significantly smaller set than $\mathcal{E}$. This can be done if the edges in the graph can be grouped as a finite number of action primitives sharing the same tunnel representation. Note that the inner optimization in the above formulation is a linear program and is guaranteed to have at least one solution as long as the feasible region of the decision variable is nonempty and bounded.
@@
The feasible region depends on the limits of the system $\boldsymbol{\phi}^{(j)}_{\text{lim}}$ and the choice of the tunnel $\tau$ which is domain-specific. However, in most cases, a straightforward design of the tunnel [45] can ensure a nonempty feasible region and a reasonably low $k_{\text{min}}$.
@@
If $\tau(e)$ is convex, then a $\boldsymbol{\phi}_{\mathbf{x}^{\prime}\mathbf{x}^{\prime\prime}}^{\sim}(t)\in\tau(e),\forall t\in[0,t_{\text{max}}]$ that solves Eq. V-B2 to global optimality can be found.
@@
Setting $\mathcal{X}^{\text{free}}=\tau(e)$ make Eq. 14d a convex domain constraint on the decision variables and Eq. V-B2 a convex program. Assuming $t_{\text{max}}$ is large enough, we know that a convex optimization can be solved to global optimality. ∎
@@
If $\boldsymbol{\hat{\phi}}_{\mathbf{x}^{\prime}\mathbf{x}^{\prime\prime}}\in\tau_{L}(e)$, then there exists a $k_{\text{min}}$-th order $\boldsymbol{\phi}_{\mathbf{x}^{\prime}\mathbf{x}^{\prime\prime}}^{*}\in\tau(e)$ that satisfies Eq. V-B2.
@@
It directly follows from the constraint satisfaction for finding $k_{\text{min}}$ in Eq. V-D1. This $k_{\text{min}}$ can be used to choose the B-spline basis for optimizing Eq. V-B2 simultaneously over path length and time to find $\boldsymbol{\phi}_{\mathbf{x}^{\prime}\mathbf{x}^{\prime\prime}}^{*}\in\tau(e)$. ∎
@@
Completeness of PINSAT: If Assumption 1 holds, then PINSAT with $k_{\text{min}}$-th order B-spline optimization is guaranteed to find a $\boldsymbol{\phi}_{\mathbf{x}^{S}\mathbf{x}^{G}}$ that satisfies Eq. III.
VI Evaluation